No apologies received.
Progress is being made on identifying the underlying cause of jobs not appearing.
Report from Computing Executive Group
Reports from units.
Hope to have this completed in a couple of weeks.
The cupboard in the AT basement server room should be rescued for future use.
Neil asked about the machines at FH on the AT wire. George mentioned that there are 2 links to FH so could move 1 at a time.
George will e-mail current users to check that they have a 2.1.x client installed.
Alastair reminded everyone that all servers should be rebooted as soon as possible, particularly ssh/login/compute servers.
The service will possibly be at risk this weekend as a result of the move.
This raised 2 meta issues:
Lammasu has now been fixed.
Neil pointed out that the fix to ensure that lpq returns information from the server has been added to the testing release and will be in the stable release this Thursday. It involves adding a line to the file /etc/cups/client.conf and has been added to the printing.h header.
Alastair noted that he has suggested that the topic of printing be discussed at CCPAG.
Topics for discussion
The next meeting will be on Wednesday 13th May in IF-G.03.
Please contact us with any comments or corrections.
Unless explicitly stated otherwise, all material is copyright The University of Edinburgh